Freddie leaves home and rides his bicycle out into the country for 2.7 hours. On his return trip along the same route, it takes him 0.9 of an hour longer. If his rate on the return trip was 3 miles per hour slower than on the trip out into the country, find the total round trip distance.
I have asked everyone in my class and no one knows how to get the answer. Instructor is unavailable until Monday. PLEASE HELP!! :) Thank you in advance

Found 2 solutions by lwsshak3, checkley79:
Answer by lwsshak3(11628)

Freddie leaves home and rides his bicycle out into the country for 2.7 hours. On his return trip along the same route, it takes him 0.9 of an hour longer. If his rate on the return trip was 3 miles per hour slower than on the trip out into the country, find the total roundtrip distance.
***
let x=rate of speed of initial trip
x-3=rate of speed of return trip
distance=travel time*speed (same for both trips)
2.7x=(2.7+.9)(x-3)
2.7x=3.6(x-3)
2.7x=3.6x-10.8
.9x=10.8
x=12
roundtrip distance=2*2.7*12=64.8 mi

Answer by checkley79(3341)

D=RT
D=R*2.7
D=(R-3)(2.7+.9)
2.7R=(R-3)*3.6
2.7R=3.6R-10.8
2.7R-3.6R=-10.8
-.9R=-10.8
R=-10.8/-.9
R=12 MPH IS THE RATE ON THE OUT TRIP.
D=12*2.7=32.4 MILES ID THE DISTACE OUT.
D=(12-3)(2.7+.9)
D=9*3.6=32.4 DITTO.
2*32.4=64.8 MILES FOR THE ROUND TRIP DIDTANCE.